if the length of a rectangle is 2cm less than 3times the width . . and the area is 30cm2 what are the dimensions of the rectangle. I think the breakdown is 3X^2 - 2X - 30 = 0 . . I am unable to factor it down . . maybe a brain lapse. . (3x+?)(x-?) ..
Found 2 solutions by lefty4ever26, mananth:
Answer by lefty4ever26(59)

Yes 3x^2 -2x-30=0 is the right equation. You have to use the Quadratic Formula to factor it though. The formula is -b plus or minus the square root of (b^2-4ac) all over 2a. It comes out with 1 (+,-) the square root of 91 all divided by 3.

Answer by mananth(16946)

I will continue from where you have left
3x^2-2x-30=0
you will have to use the quadratic equation
discriminant = 364
x1= (2+sqrt(364))/6}}}
x1= 3.51
x2 will be negative so ignore.
width = 3.51 cm
length = 3x-2
3*3.51-2
=10.53-2
=8.53
..
Check
8.53*3.51= 29.94
...
m.ananth@hotmail.ca